# Inflation, the Hubble Tension and Early Dark Energy: an alternative overview

### v1

@@ III.3 Datasets
The datasets involved in my analysis are:
  • Measurements of CMB temperature anisotropy and polarization power spectra, their cross-spectra, and the lensing power spectrum reconstructed from the temperature 4-point correlation function, all from the Planck 2018 legacy data release, using the plik likelihoods Aghanim et al. (2020b, c, d);
  • B-modes CMB polarization data released by the BICEP/Keck Collaboration Ade et al. (2021);
@@ IV Early-time new physics and Inflation
In this section, I study the implications for inflation resulting from early-time new physics parametrized by imposing $N_{\rm eff}>3.044$. I divide the section into two subsections: in subsection IV.1, I discuss the implications for generic single-field models of inflation where the inflationary potential is parameterized in terms of the slow-roll parameters. Conversely, in subsection IV.2, I will focus on the implications of one particular benchmark scenario: the model proposed by A. Starobinsky Starobinsky (1980).
@@ IV.1 Implications for single field inflation
$n_{s}-1=2\eta-6\epsilon,\quad n_{\rm t}=-2\epsilon=-r/8.$ (10)
@@
In the context of single-field slow-roll inflation, the values of the spectral index and tensor amplitude can be directly correlated to the values of derivatives of the inflationary potential through the relationships indicated in Eq. (10). Therefore, another point I would like to briefly discuss concerns the possible implications of the Hubble tension for inflationary models. Keeping the discussion as general as possible, in Figure 2, I depict the constraints on the two slow-roll parameters, $\epsilon$ and $\eta$, as well as their correlation with the value of $r_{\rm{drag}}$. As highlighted in the figure, the parameter $\epsilon$ (which determines the amplitude of tensor perturbations) is not sensitive to the value of the sound horizon. For all cases considered, we obtain a limit of $\epsilon\lesssim 0.002$. Conversely, the significant shift in the value of $n_{s}$ is translated into a variation in the predicted value for the second slow-roll parameter, $\eta$. Within the standard cosmological model, $\eta$ is measured to be $\eta=-0.0130^{+0.0024}_{-0.0029}$, namely $\eta<0$ and such that $1\gg|\eta|\gg\epsilon$. Being linked to derivatives of the inflationary potential, such a hierarchy poses a minor fine-tuning problem since there exists no fundamental reason or symmetry in nature that can justify it. In fact, it simply stands as a prediction of current data imposing strong constraints on the inflationary models that can be deemed suitable to explain current observations. As a matter of fact, a few well-motivated inflation models that are able to predict such a hierarchy have become along the years benchmark scenarios to gauge future CMB experiments. This is, for instance, the case of the model proposed by Alexei Starobinsky in Ref. Starobinsky (1980), which is able to naturally explain such a hierarchy between the inflationary slow-roll parameters. That said, observing Figure 2 again, we can notice how this hierarchy is called into question when considering the possibility of introducing new physical mechanisms that could resolve the Hubble tension. As soon as the value of the sound horizon gets smaller (and $H_{0}$ higher), the parameter $\eta$ shifts towards $\eta\to 0$, allowing in some cases values close to 0 or even crossing the positive region. While this could reduce the fine-tuning between the two slow-roll parameters, it has important implications in terms of model selection, as discussed in the following subsection.
@@ IV.2 Implications for Starobinsky inflation
where $a_{0}H_{0}\simeq 2.248\times 10^{-4}\rm{Mpc}^{-1}$ represents the inverse of the comoving horizon size in the current Universe, $H_{*}$ is the value of the Hubble parameter at the horizon exit, $H_{\rm end}$ is the Hubble parameter at the end of inflation, $\Lambda$ is the energy scale of inflation and the subscript "eq" denotes quantities evaluated at matter-radiation equality. Approximating $H_{\rm end}\simeq H_{*}$ and recalling that in single field models of inflation the energy-scale can be related to the amplitude of tensor perturbations by $\Lambda\simeq 3.3\times r^{1/4}\times 10^{16}\rm{GeV}$, one can get a rough idea of the value of $\mathcal{N}$ at $k=k_{*}=0.05$ Mpc-1. In particular, fixing a $\Lambda$CDM-like history after inflation, the typical range allowed for $\mathcal{N}$ within single-field models falls in the range $\mathcal{N}\in[55,65]$  Liddle and Leach (2003) [^10] It is theoretically possible to achieve a greater number of $e$-folds of expansion by reducing the energy-density during inflation. This can result in an extreme upper bound $\mathcal{N}\lesssim 100$, see Ref. Liddle and Leach (2003).. Plugging these values into Eq. (15) makes it easy to see that the value obtained for the inflationary parameters perfectly aligns with those measured by the Planck collaboration (as well as with the upper bound on $r$ obtained by BK18 data) when a $\Lambda$CDM model of cosmology is assumed.
@@ VI.4 Early Dark Energy and Starobinsky Inflation
Before proceeding further, I would like to emphasize that this question has been already partially discussed in the literature, although with different levels of detail. In particular, in Refs. Jiang et al. (2022); Ye et al. (2022); Ye and Piao (2022), it has been argued that reducing the Hubble constant discrepancy within EDE requires a significant change in $n_{s}$ that is in overall disagreement with the theoretical predictions of the aforementioned inflation model. That being said, it is important to note that these results were obtained without assuming any specific inflation model, but only comparing the results of inflationary parameters with those predicted by Eqs. (15) afterwards. Here, I want to follow a different approach and consider both these models together from the beginning of the analysis. In other words, I assume the theoretical relations between $n_{s}$, $r$, and $\mathcal{N}$ predicted by the Starobinsky model from scratch, thereby reducing the number of free parameters. For EDE, I consider the same numerical implementation discussed in  subsection VI.1. As a result, the cosmological model I am considering in this subsection has 9 free parameters: $\omega_{\rm b}\equiv\Omega_{\rm b}h^{2}$, $\omega_{\rm c}\equiv\Omega_{\rm c}h^{2}$, $\theta_{*}$, $\tau$, $\log(10^{10}A_{\rm s})$, $\mathcal{N}$, $f_{\rm EDE}(z_{c})$, $z_{c}$, $\theta_{i}$}. I analyze this model in light of observations from Planck+BK18+BAO and Planck+BK18+BAO+$H_{0}$.
@@
When considering a prior on $H_{0}$ from SH0ES, one expects the contribution of models with larger $f_{\text{EDE}}$ to become much more significant, as these models can increase $H_{0}$ and better explain data. This is what I observe in Figure 9: assuming a prior on $H_{0}=73\pm 1$ km/s/Mpc, the black contours shift significantly in the direction of a non-negligible fraction of Early Dark Energy and a lower value of the sound horizon. This implies higher values of $n_{s}$ and, within the Starobinsky model, higher values of $\mathcal{N}$. Indeed, as evident from Figure 10, values of the number of e-folds in the range $\mathcal{N}\in[50,70]$ are now excluded at more than 3 standard deviations. This unequivocally confirms that the fraction of Early Dark Energy allowed to maintain agreement with the theoretical predictions of the Starobinsky model cannot be as large as required to resolve the Hubble tension. Therefore, the conclusion I draw is that resolving the Hubble tension in the context of Early Dark Energy excludes Starobinsky Inflation at more than $3\sigma$ as the mechanism describing inflation.

### v2

@@ III.3 Datasets
The datasets involved in my analysis are [^9] Concerning the choice of likelihoods employed in my study, I conservatively used the very same combinations of observables analyzed in the Planck 2018 paper on inflation Akrami et al. (2020a). Namely the same Planck 2018 CMB likelihoods (including the lensing potential), B-mode polarization measurements, and baryon acoustic oscillations.:
  • Measurements of the Planck CMB temperature anisotropy and polarization power spectra, their cross-spectra, and the lensing power spectrum. All the Planck CMB likelihoods employed in this work are listed below:
    • (i)
      Measurements of the power spectra of temperature and polarization anisotropies, $C_{\ell}^{TT}$, $C_{\ell}^{TE}$, and $C_{\ell}^{EE}$, at small scales ($\ell>30$), obtained by the Planck plik likelihood Aghanim et al. (2020b, c);
    • (ii)
      Measurements of the spectrum of temperature anisotropies, $C_{\ell}^{TT}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck Commander likelihood Aghanim et al. (2020b, c);
    • (iii)
      Measurements of the spectrum of E-mode polarization, $C_{\ell}^{EE}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck SimAll likelihood Aghanim et al. (2020b, c);
    • (iv)
      Reconstruction of the spectrum of the lensing potential, obtained by the Planck plik likelihood Aghanim et al. (2020d).
  • B-modes CMB polarization data released by the BICEP/Keck Collaboration Ade et al. (2021).
@@ IV Early-time new physics and Inflation
In this section, I study the implications for inflation resulting from early-time new physics parametrized by imposing $N_{\rm eff}>3.044$. I divide the section into two subsections: in subsection IV.1, I discuss the implications for generic single-field models of inflation where the inflationary potential is parameterized in terms of the slow-roll parameters. Conversely, in Figure 3, I will focus on the implications of one particular benchmark scenario: the model proposed by A. Starobinsky Starobinsky (1980).
@@ IV.1 Implications for single field inflation
$n_{s}-1=2\eta-6\epsilon,\quad n_{\rm t}=-2\epsilon=-r/8,$ (10)
where, as usual, the slow-roll parameters $\epsilon$ and $\eta$ are calculated for specific field values $\phi_{*}$ corresponding to when perturbations cross the horizon during inflation.
@@
In the context of single-field slow-roll inflation, the values of the spectral index and tensor amplitude can be directly correlated to the values of derivatives of the inflationary potential through the relationships indicated in Eq. (10). Therefore, another point I would like to briefly discuss concerns the possible implications of the Hubble tension for inflationary models. Keeping the discussion as general as possible, in Figure 2, I depict the constraints on the two slow-roll parameters, $\epsilon$ and $\eta$, as well as their correlation with the value of $r_{\rm{drag}}$. As highlighted in the figure, the parameter $\epsilon$ is not sensitive to the value of the sound horizon. This parameter is directly related to the amplitude of tensor perturbations by the slow roll relation $r=16\epsilon$. B-mode polarization measurements from BICEP/Keck likelihood set very stringent constraints on the tensor amplitude $r$, and these constraints are essentially unchanged with the cosmological model [^11] The limit on the tensor amplitude is primarily due to measurements of the BB angular power spectrum at large angular scales ($\ell\lesssim 100$), where the lensing contribution in the BB spectrum is subdominant compared to the primordial tensor counterpart. In this range of the spectrum, the only significant parameter to gauge the amplitude of the signal is the tensor-to-scalar ratio $r$, significantly reducing the dependence of the constraints on the background cosmology.. As a result, for all cases considered, I obtain a very tight limit of $\epsilon\lesssim 0.002$, making its contribution in Eq. (10) subdominant compared to $\eta$ [^12] I note that in the minimal theoretical setup described by single field slow-roll inflation, it must be $\epsilon>0$ to preserve the Null Energy Condition. Therefore, this parameter is bounded in the range $\epsilon\in[0,0.002]$.. Therefore, the significant shift in the value of $n_{s}$ is recast into a variation in the inferred value for the second slow-roll parameter, $\eta$. Within the standard cosmological model, $\eta$ is measured to be $\eta=-0.0130^{+0.0024}_{-0.0029}$, namely $\eta<0$ and such that $1\gg|\eta|\gg\epsilon$. Being linked to derivatives of the inflationary potential, such a hierarchy poses a minor fine-tuning problem since there exists no fundamental reason or symmetry in nature that can justify it. In fact, it simply stands as a prediction of current data imposing strong constraints on the inflationary models that can be deemed suitable to explain current observations. As a matter of fact, a few well-motivated inflation models that are able to predict such a hierarchy have become along the years benchmark scenarios to gauge future CMB experiments. This is, for instance, the case of the model proposed by Alexei Starobinsky in Ref. Starobinsky (1980), which is able to naturally explain such a hierarchy between the inflationary slow-roll parameters. That said, observing Figure 2 again, we can notice how this hierarchy is called into question when considering the possibility of introducing new physical mechanisms that could resolve the Hubble tension. As soon as the value of the sound horizon gets smaller (and $H_{0}$ higher), the parameter $\eta$ shifts towards $\eta\to 0$, allowing in some cases values close to 0 or even crossing the positive region [^13] To provide concrete numbers quantifying this shift, I observe that within the $\Lambda$CDM model, $\eta=0$ is ruled out at $5.4\sigma$, whereas for $\Delta N_{\rm eff}\gtrsim 0.6$, it is well allowed within the 68% confidence level contours.. While this could reduce the fine-tuning between the two slow-roll parameters, it has important implications in terms of model selection. Referring back to Eq. (9) we see that determining whether $\eta$ is positive or negative is equivalent to determining the sign of the second derivative of the inflationary potential $V^{\prime\prime}(\phi)$, and thus whether the inflationary potential is concave or convex. Maintaining a standard cosmological model, the various datasets analyzed in this study produce probability contours in the concave region ($\eta<0$), while altering the cosmological model the results gradually shift towards the region of convex potentials ($\eta>0$), thereby challenging some well-established benchmark models, as discussed in the next section.
@@ IV.2 Implications for Starobinsky inflation
where $a_{0}H_{0}\simeq 2.248\times 10^{-4}\rm{Mpc}^{-1}$ represents the inverse of the comoving horizon size in the current Universe, $H_{*}$ is the value of the Hubble parameter at the horizon exit, $H_{\rm end}$ is the Hubble parameter at the end of inflation, $\Lambda$ is the energy scale of inflation and the subscript "eq" denotes quantities evaluated at matter-radiation equality. Approximating $H_{\rm end}\simeq H_{*}$ and recalling that in single field models of inflation the energy-scale can be related to the amplitude of tensor perturbations by $\Lambda\simeq 3.3\times r^{1/4}\times 10^{16}\rm{GeV}$, one can get a rough idea of the value of $\mathcal{N}$ at $k=k_{*}=0.05$ Mpc-1. In particular, fixing a $\Lambda$CDM-like history after inflation, the typical range allowed for $\mathcal{N}$ within single-field models falls in the range $\mathcal{N}\in[55,65]$  Liddle and Leach (2003). Plugging these values into Eq. (15) makes it easy to see that the value obtained for the inflationary parameters perfectly aligns with those measured by the Planck collaboration (as well as with the upper bound on $r$ obtained by BK18 data) when a $\Lambda$CDM model of cosmology is assumed.
@@ VI.4 Early Dark Energy and Starobinsky Inflation
Before proceeding further, I would like to emphasize that this question has been already partially discussed in the literature, although with different levels of detail. In particular, in Refs. Jiang et al. (2022); Ye et al. (2022); Ye and Piao (2022); Fu and Wang (2024), it has been argued that reducing the Hubble constant discrepancy within EDE requires a significant change in $n_{s}$ that is in overall disagreement with the theoretical predictions of the aforementioned inflation model. That being said, it is important to note that these results were obtained without assuming any specific inflation model, but only comparing the results of inflationary parameters with those predicted by Eqs. (15) afterwards. Here, I want to follow a different approach and consider both these models together from the beginning of the analysis. In other words, I assume the theoretical relations between $n_{s}$, $r$, and $\mathcal{N}$ predicted by the Starobinsky model from scratch, thereby reducing the number of free parameters. For EDE, I consider the same numerical implementation discussed in  subsection VI.1. As a result, the cosmological model I am considering in this subsection has 9 free parameters: $\omega_{\rm b}\equiv\Omega_{\rm b}h^{2}$, $\omega_{\rm c}\equiv\Omega_{\rm c}h^{2}$, $\theta_{*}$, $\tau$, $\log(10^{10}A_{\rm s})$, $\mathcal{N}$, $f_{\rm EDE}(z_{c})$, $z_{c}$, $\theta_{i}$}. I analyze this model in light of observations from Planck+BK18+BAO and Planck+BK18+BAO+$H_{0}$.
@@
When considering a prior on $H_{0}$ from SH0ES, one expects the contribution of models with larger $f_{\text{EDE}}$ to become much more significant, as these models can increase $H_{0}$ and better explain data. This is what I observe in Figure 9: assuming a prior on $H_{0}=73\pm 1$ km/s/Mpc, the black contours shift significantly in the direction of a non-negligible fraction of Early Dark Energy and a lower value of the sound horizon. This implies higher values of $n_{s}$ and, within the Starobinsky model, higher values of $\mathcal{N}$. Indeed, as evident from Figure 10, values of the number of e-folds in the range $\mathcal{N}\in[50,70]$ are now excluded at more than 3 standard deviations. In this regard, it is worth noting that the exact number of e-folds of inflation $\mathcal{N}$, generally depends on the model itself and the details of the subsequent post-inflationary evolution through the reheating temperature Lidsey et al. (1997); Liddle and Leach (2003). Although under some circumstances it is theoretically possible to achieve models with a greater number of e-folds [^16] In general, to achieve a greater number of $e$-folds of expansion, one needs to reduce the energy density during inflation or significantly lower the energy scale at the end of inflation. In these non-standard scenarios, it is possible to achieve an extreme upper bound of $\mathcal{N}\lesssim 100$; see Ref. Liddle and Leach (2003)., these realizations are typically not applicable to the minimal $R^{2}$-inflation considered in this work. In realistic realizations of reheating in Starobinsky inflation Gorbunov and Panin (2011); Bezrukov and Gorbunov (2012), the inflaton coupling to all fields is Planck scale suppressed, and reheating mainly occurs via its decays into the Standard Model Higgs bosons (which immediately rescatter into Standard Model particles) since the inflaton couplings to all other fields are further suppressed due to conformal symmetry Starobinsky (1980); Vilenkin (1985). Consequently, the reheating temperature in Starobinsky inflation is usually lower than in other benchmark scenarios and can be estimated to be of the order of $T_{\text{rh}}\sim 10^{9}$ GeV. Therefore even in the most optimistic – yet realistic – setup, one gets a number of e-folds not greater than $\mathcal{N}\sim 55-60$, see e.g, Ref Bezrukov and Gorbunov (2012) (or Ref. Akrami et al. (2020a)). This unequivocally confirms that the fraction of Early Dark Energy allowed to maintain agreement with the theoretical predictions of the Starobinsky model cannot be as large as required to resolve the Hubble tension. Therefore, the conclusion I draw is that resolving the Hubble tension in the context of Early Dark Energy excludes Starobinsky Inflation at more than $3\sigma$ as the mechanism describing inflation.